Martin Office Supplies paid a $5 dividend last year. The dividend is expected to grow at a constant rate of 10 percent over the next four years. The required rate of return is 17 percent (this will also serve as the discount rate in this problem). Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods.

a. Compute the anticipated value of the dividends for the next four years. (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

b. Calculate the present value of each of the anticipated dividends at a discount rate of 17 percent. (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

D1 =

D2 =

D3 =

D4=

Total$0.00

c. Compute the price of the stock at the end of the fourth year (P4). (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

d. Calculate the present value of the year 4 stock price at a discount rate of 17 percent. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

e. Compute the current value of the stock. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

f. Use the formula given below to show that it will provide approximately the same answer as part e. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

P0 =

D1

Keg

g. If current EPS were equal to $7.00 and the P/E ratio is 13% higher than the industry average of 10, what would the stock price be? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

h. By what dollar amount is the stock price in part g different from the stock price in part f? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

i. With regard to the stock price in part f, indicate which direction it would move if:

(1) D1 increases

(2) Ke increases

(3) g increases
Answer #1

a.Expected Dividend = Prior Dividend(1+Growth rate)

D1 = 5(1.10) = $5.5

D2 = 5(1.10)2 = $6.05

D3 = 5(1.10)3 = $6.655

D4 = 5(1.1)^4 = $7.3205

b.Present value = 5.5/1.17 + 6.05/(1.17)2 + 6.66/(1.17)3 + 7.32/(1.17)4

= $17.19

c.Price = Dividend in Year 5/(Discount rate – growth rate)

= 7.32(1.10)/(17%-10%)

= $115.03

Present value = 115.03/(1.17)4

= $61.39

Current value = 17.19+61.39

= $78.58

e.Stock Price = 5.5/(17%-10%)

= $78.57

g.Stock Price = EPS*P/E Ratio

= 7*10*1.13

= $79.10

h.Differebce = $0.53

Price will increase

Price will reduce

Price will increase